Difference frequency lasers mix two beams of radiation, usually in the visible, at frequencies ν1 and ν2 to generate infrared radiation at the frequency ν3=|ν2−ν1|.

Part A Find the wavelength in microns of the difference frequency laser radiation generated by mixing beams with wavelengths 612 nm and 438 nm .

Express your answer using three significant figures. λ3 =
Answer #1

Given that λ1 = 612 nm , λ2 = 438 nm

ν3=|ν2−ν1|

We know that c = vλ

v = c/λ

v and λ are inversely related.

Therefore, 1/λ3 = |1/λ2 - 1/λ1|

  1/λ3 =  |λ1- λ2/ λ1.λ2|

  λ3 = λ1.λ2 /  |λ1- λ2|

= (612 nm. 438 nm) / |612 nm-438 nm|

= 1540.5 nm

Theerfore,  λ3 = 1540.5 nm